The following balanced reaction was performed . If 558g of Mg was reacted, what is the total energy released from the reaction? MM(Mg)= 24.31g/mol

2(Mg)s +O2(g)= 2Mgo + 1203KJ/mol
1
Expert's answer
2020-12-06T10:34:02-0500

n(Mg)=m/M=558g / 24.31 g/mol = 22.954 mol

According to reaction:

"\Sigma"Q = Qreaction/2 * n(Mg) = 1203/2 (kJ/mol) * 22.954 = 13806.5 kJ = 13.8 MJ